y=5x-2
-3x+6y=-12

Answer by Boreal(15235)

y=5x-2. Substitute into the second
-3x+6(5x-2)=-12
-3x+30x-12=-12
The -12 cancel
27x=0
x=0
y=-2
check in the second equation
0+6(-2)=-12
(0,-2)